# An Investigation into the Performances of the State-of-the-art Machine Learning Approaches for Various Cyber-attack Detection: A Survey

### v1

@@ Abstract
To secure computers and information systems from attackers taking advantage of vulnerabilities in the system to commit cybercrime, several methods have been proposed for real-time detection of vulnerabilities to improve security around information systems. Of all the proposed methods, machine learning had been the most effective method in securing a system with capabilities ranging from early detection of software vulnerabilities to real-time detection of ongoing compromise in a system. As there are different types of cyberattacks, each of the existing state-of-the-art machine learning models depends on different algorithms for training which also impact their suitability for detection of a particular type of cyberattack. In this research, we analyzed each of the current state-of-the-art machine learning models for different types of cyberattack detection from the past 10 years with a major emphasis on the most recent works for comparative study to identify the knowledge gap where work is still needed to be done with regard to detection of each category of cyberattack.
@@ II Background Study
II-A Distributed Denial of Service (DDoS) Attack
In a distributed denial of service attack, the resources of a targeted system are maliciously flooded from multiple systems at a time to disrupt the normal traffic flow of the targeted server, service, or network, in most cases the traffic is synthetically generated by the attacker to maliciously overwhelm the system which eventually leads to denial of service as the system is unable to deliver service to legitimate user. It uses brute force attack [75] which can be triggered through Botnet when any of the devices of the network environment are actively infected with malware. DDoS attacks can be classified into three main categories which could be (i) Traffic/fragmentation attack, (ii) Bandwidth/Volume attack and (iii) Application attack [75] depending on the nature, severity of the attack, or form of the attack.
II-B Malware Attack
Malware attack is a type of cyberattack that are mostly introduced through phishing, social engineering, and downloads in which unauthorized actions by malicious software commonly known as a virus are executed on the victim’s system. Each malicious software has different malicious activities ranging from stealing sensitive data, launching DDoS attacks [28], conducting ransomware, or pushing unwanted adverts in the case of adware attacks, these specialized activities determine the type of malware attack which in some cases might be command and control, spyware, ransomware, adware and so on.
II-C Man in the Middle (MITM) Attack
Man-in-the-middle[a] (MITM) attack is a type of cyberattack where an attacker secretly relays or alters communications between two parties. By intercepting and relaying the messages back to the intending receiver, the attacker makes the sender and the receiver believe they are directly communicating with each other. It is a sort of eavesdropping where the attacker connects with the victim’s network to relay incoming and outgoing messages which could be altered by the attacker before relaying back to the receiver [50]. Man-in-the-browser is the most common MITM attack where the attacker could inject malicious proxy malware into the victim’s device through browser infection. MITN attack is carried out by data interception involving the interception of a data transfer between client and server and the actual decryption of the messages. Malware that is involved in this type of attack is mostly introduced to the victim’s system by phishing email where the goal is to steal sensitive and other personal information from the victim.
II-D Drive-by Attack
A drive-by attack is the malicious infection of a computer with malware when a victim visits an infected website. Merely visiting an infected website is enough to get the malware downloaded and running on a system unknown to the victim. This type of attack exploits vulnerabilities in the victim’s browser application to successfully infect the computer. Still, they become obsolete when a security patch for the browser application is released. The drive-by attack uses different malware distribution techniques, sophistication, and attack intensity which can go unnoticed for a long time [78] and with the ability to cause significant damage to a system, as the process involves attacking a legitimate website by injecting malicious code into the pages to compromise it so that the injected code could be loaded into the victim’s system through his browser whenever a user browses the compromised legitimate website thereby initializing the drive-by attack
II-E SQL Injection Attack
SQL injection is an application layer attack through which hackers steal sensitive data by inserting malicious SQL statements into the input field of an application for execution such as the input of SQL query to dump the database contents to the attacker. In this type of attack, the aim of the attacker is the usage of malicious SQL code for backend database manipulation to access information that is not supposed to be displayed [22] such as private customer details, company data, user’s list, and so on.
II-F Phishing Attack
Phishing is a type of cybercrime in which an individual is lured to divulging sensitive information details through text message, email, or phone conversation by someone posing either as a legitimate institution or a member of a legitimate institution, some of these commonly requested sensitive details which are social security number, password, credit, and banking card details etc are later used to access more sensitive information [56] for different type of cybercrime which often results in financial loss or identity theft as about 76% of the phishing attacks were credential-harvesting in 2022 according to Digital Information world. A California teenager was able to get sensitive information to access credit card details and withdraw money from his victim’s account through his fake ”America Online” website which resulted in the first lawsuit filed in 2004. Efficient phishing detection has been challenging as attackers continue to advance their tactics as technologies evolve. To defraud personnel, all an attacker needs to do is simply clone a legitimate website to create a new website (SCAM Website) which is then used to defraud computer users.
III Category of Current State of the art Phishing Detection model
III-A Bayesian-Based-Classifier
@@
III-B Non-Bayesian Based Classifier
III-B1 Decision Tree
@@ III-B5 K-Nearest Neighbor (KNN)
III-B6 Support Vector Machine (SVM)
@@
III-C Deep Learning Based Classifier
III-C1 Convolutional Neural Network (CNN)
@@
III-C2 Recurrent Neural Network (RNN)
@@
III-C3 Long Short-Term Memory (LSTM)
@@ IV Literature Review
Mayank Agarwal et al.[3] proposed IDS that adheres to the 802.11 standard which does not require any protocol modification to detect the existence of flooding-based DoS attacks in a Wi-Fi network and to quickly recover from the attack. The proposed intrusion detection system is independent of client software and only requires hardware that is capable of sniffing the wireless data to ensure cost-effectiveness.
Figure 1: MITM Attack detection rate, accuracy, and differences from different classifiers from Mayank Agarwal et al [3]
For man-in-the-middle Attack Detection 1, Naive Bayes has the weakest overall performance with an accuracy of 68%[3], while another probabilistic classifier in Bayes Net has significant improvement in performance over Naive Bayes with accuracy of 95% and a detection rate of 88% while SVM algorithm has the best performance in terms of accuracy with accuracy rate of 98.7% but with abysmal detection rate of 57.8% which is extremely poor for an intrusion detection system.
Muhanna Saed and Ahamed Aljuhani [73] proposed a set of machine learning techniques both to detect and identify Man-In-The-Middle attacks on a wireless communication network. In addition, validation and evaluation were based on performance metrics, as well as performance comparison with other machine learning-based Man-In-The-Middle attack detection methods. By training traditional machine learning-based and deep learning-based models with a set of data that represents the transmission of data over a wireless network for MITM attack detection, the deep learning-based model which was trained with Long short-term memory (LSTM) network as a recurrent neural network (RNN) as an accuracy of 92%, Support vector machine (SVM) having an accuracy of 85%, while random forest with accuracy of 94%.
Ann Zeky et al [49] proposed extraction based naive bayes model for phishing detection with emphasis on the extraction of relevant features like unusual characters, spelling mistakes, domain names and URL analyzation with substantial success albeit with an imbalance dataset and susceptibility to bayesian poisoning.
Patricia Iglesias [39] used an artificial neural network for the detection of drive-by attacks on polyglot payloads in the image, however, the results obtained are only limited to the successful detection of stego images by the use of LSB and F5 steganographic methods. Hence, they propose deep learning techniques through the use of convolutional Neural Network as a suitable method of detection in a situation where both the malicious content and the images are delivered by Spatial and Steganography algorithms. The proposed CNN model was evaluated with benchmark image databases along with collections of JavaScript exploits which yielded an AUC validation score of 99.75% and an accuracy of 98.61%.
mahdi bahaghighat et al [18] proposed phishing detection method based on Logistic Regression, K-Nearest Neighbors, Naive Bayes, Random Forest, Support Vector Machine, and Extreme Gradient Boosting (XGBoost) algorithm relying solely on the attribute feature of the webpage URL. The experiment resulted in the outperformance of other algorithms by XGBoost, Random Forest, KNN with accuracy of 99.2%, 98.1% and 98.3% respectively while naive bayes has the worst performance with accuracy of 93% for phishing detection of all the algorithm echoing result obtained by Kamal Omari [58] on the performance of naive bayes algorithm for phishing detection.
Morufu Olalere et al. [57] proposed a naïve Bayes model for effective categorization and detection of SQL injection attacks, they obtained 98% and 99% for detection and categorization respectively by validating the proposed model with stratified cross validation having 1-10 random seeds which is a significant leap when compared with previous studies and performance, although we don’t know the actual distribution of different categories of SQL injection attack type in the dataset, and also how the proposed Naive Bayes model will perform on KAGGLE SQL injection attack dataset.
While assuming the absence of a single solution to detect most phishing attacks, and to evaluate performances of Bayesian classier for phishing detection tasks based on different feature selection techniques. Twana Mustafa and Murat Karabatak [53] developed 6 bayesian based models in which each model involves a single feature selection technique chosen from individual FS, forward FS, Backward FS, Plus-I takeaway-r FS, AR1, and All to compare the performance of bayesian classifier for phishing detection task based on different feature selection techniques. The experiment resulted in the Bayesian model with Plus-I takeaway-r feature selection having the best performance with an accuracy of 93.39% while the Bayesian classifier with individual feature selection technique has the least performance with an accuracy of 92.05%, and so concluded that feature selection has a direct impact on classifier accuracy.
To investigate the performance of different classifiers in detecting SQL Injection attacks, Prince Roy et al. [69] trained multiple models with Logistic Regression, AdaBoost (Adaptive Boosting), Random Forest, Naive Bayes, and XGBoost (Extreme Gradient Boosting) Classifier on Kaggle SQL Injection Dataset, and concluded with Naive Bayes as the best classifier for SQL injection attack detection with an accuracy of 98.33% echoing similar conclusion from Morufu Olalere [57]. Logistic Regression provided 92.73% accuracy, Adaboost provided 90.35% accuracy, XGBoost provided 89.64%, and Random Forest provided 92.14% accuracy.
TABLE I: Comparison Analysis of Methodologies for Detection of SQL Injection Attack
Classifier
Authors
Mean Score
Naive Bayes
[69], [87], [23], [46], [61], [77], [8]
90.4
SVM
[87], [23], [31], [46], [77], [34],[8]
87.63
Random Forest
[69], [87], [77], [34], [82], [9], [8]
93.71
Logistic Regression
[69], [46], [34], [16], [9], [8], [19]
89.68
KNN
[77], [9], [8], [2], [32], [19], [41]
87.2
Decision Tree
[34], [23], [82], [8], [2], [32], [40]
90.04
TABLE II: Comparison Analysis of Methodologies for Detection of DDoS Attack
Classifier
Authors
Mean Score
SVM
[81], [48], [33], [14], [62], [83], [51]
90.0
Naive Bayes
[81], [74], [33], [4], [67], [91], [14]
84.6
Random Forest
[7], [74], [33], [27], [4], [67], [21]
93.34
Decision Tress
[7], [81], [33], [27], [14], [83], [64]
96
XGBoost
[7], [27], [51], [21], [70], [24]
96.2
KNN
[81], [27], [14], [25], [90], [54], [64]
96.5
TABLE III: Comparison Analysis of Methodologies for Detection of Phishing Attack
Classifier
Authors
Mean Score
Naive Bayes
[18], [65], [89], [43], [58], [5], [6], [1], [59], [84], [68], [76], [72], [13], [45]
80.431
SVM
[18], [65], [43], [58], [12], [5], [66], [15], [60], [85], [44], [13]
89.429
Random Forest
[18], [65], [89], [43], [58], [11], [12], [5], [60], [10], [66], [71],[44]
97.065
Decision Tree
[65], [43], [58], [12], [5], [66], [60], [52], [44], [59], [84], [13]
95.248
Logistic Regression
[18], [65], [58], [15], [60], [80], [71], [44], [47], [1], [59]
92.589
KNN
[18], [65], [43], [58], [5], [10], [15], [60], [80], [71], [20], [44], [84], [13]
90.479
TABLE IV: Limitations of current state-of-the-art Approaches for cyberattack detection
Author
Dataset
Research Summary
Method/Algorithm
Limitation
Mayank Agarwal et al.[3]
Internally generated dataset
Proposal of IDS to detect the existence of flooding-based DoS attacks in a Wi-Fi network
Naive Bayes, Bayes Net, Ridor, ADTree
1. The dataset used is not publicly available
2. An unknown number of records in the generated dataset
Ali Bin et al. (2022) [79]
Kali Linux distribution having 3855 attacks
evaluation of machine learning classifiers for the detection of man-in-the-middle attacks
Decision Tree
Naive Bayes
K-Nearest Neighbor(KNN)
Artificial Neural Network
XGBoost
2. The result obtained is limited to only one dataset
Muhanna Saed and Ahamed Aljuhani (2022) [73]
Wi-Fi network benchmark dataset
Proposal machine learning techniques both to detect and identify Man-In-The-Middle attacks on a wireless communication network
LSTM, Support Vector Machine, Random Forest
1. proposed model likely to be bias due to imbalance dataset
Kamal Omari (2023) [58]
UCI phishing domains dataset
an investigation into the performances and efficiency of Logistic Regression, KNN, SVM, Naive Bayes, Decision Tree, Random Forest, and Gradient Boosting for phishing detection task
naive bayes
KNN
SVM
Random Forest
Gradient Boost
Decision Tree
Logistic Regression
Prince Roy et al. (2022) [79]
Kaggle SQL Injection attack dataset
evaluation of machine learning classifiers for the detection of SQL Injection attacks
Logistic Regression
Naive Bayes
XGBoost
Random Forest
Adaboost
Ann Zeky et al (2023) [49]
Internally generated dataset
proposal of extraction based naive Bayes robust model for phishing detection with emphasis on the combination of webpage content and URL feature analysis
naive Bayes
URL analyzation
webpage content extraction
Patricia Iglesias et al. [39]
Coco and ILSVR dataset
proposal of deep learning based CNN model for detecting drive-by road attack
Convolutional Neural Network
limited to drive-by road attack detection in stego images
mahdi bahaghighat (2023) [18]
Grega Vrbančič phishing dataset
performance comparison of phishing detection method based on several six different algorithm
naive bayes
KNN
SVM
Random Forest
Gradient Boost
Logistic Regression
Nishitha U et al (2023) [55]
Kaggle Phishing Dataset
performance comparison of machine learning and deep learning based algorithm for phishing detection
CNN
RNN
KNN
Random Forest
Decision Tree
Logistic Regression
2. imbalance in the dataset will lead to bias
sole reliance on URL feature
Prince Roy et al. (2022) [79]
Kaggle SQL Injection attack dataset
evaluation of machine learning classifiers for the detection of SQL Injection attacks
Logistic Regression
Naive Bayes
XGBoost
Random Forest
Adaboost
Twana Mustafa and Murat Karabatak (2023) [53]
UCL Phishing Dataset
Performance Comparison of different Bayesian classifier based on different Feature Selection Algorithm
naive bayes
individual FS
forward FS
backward FS
Plus-I takeaway-r FS
AR1 FS
no hint on why Plus-I takeaway-r FS works better best for naive bayes
Morufu Olalere
dataset was internally generated from dmoz
proposal of Naive Bayes model for detecting SQL injection attack
Naive Bayes
SQL syntax distribution in the dataset was not known, It is difficult to know how the proposed model will perform on Kaggle SQL injection dataset
Jaya T et al (2023) [42]
UK-2011, SpamBase, and Spam Assassin datasets
usage of frequency weightage of the words for unsupervised clustering of mail into spam and ham messages
naive bayes
random forest
logistic regression
random tree
LTSM
Santhosh Raminedi et al (2023) [65]
Internally collected phishing dataset
evaluation of several machine learning and deep learning based algorithms for phishing detection using URL features
ANN
SVM
KNN and Naive Bayes
Random Forest
Decision Tree
Logistic Regression
2. sole reliance on URL feature
V Analysis and Discussion
@@ 3. Limitation of Current Approach to SQL Injection Attack Detection
VI Conclusion and Research Direction

### v2

@@ Abstract
In this research, we analyzed the suitability of each of the current state-of-the-art machine learning models for various cyberattack detection from the past 5 years with a major emphasis on the most recent works for comparative study to identify the knowledge gap where work is still needed to be done with regard to detection of each category of cyberattack. We also reviewed the suitability, effeciency and limitations of recent research on state-of-the-art classifiers and novel frameworks in the detection of differnet cyberattacks. Our result shows the need for; further research and exploration on machine learning approach for the detection of drive-by download attacks, an investigation into the mix performance of Naive Bayes to identify possible research direction on improvement to existing state-of-the-art Naive Bayes classifier, we also identify that current machine learning approach to the detection of SQLi attack cannot detect an already compromised database with SQLi attack signifying another possible future research direction.
@@ I Introduction
TABLE I: Limitations of current state-of-the-art Approaches for cyberattack detection
Author Dataset Research Summary Method/Algorithm Limitation
Mayank Agarwal et al.[4] Internally generated dataset Proposal of IDS to detect the existence of flooding-based DoS attacks in a Wi-Fi network Naive Bayes, Bayes Net, Ridor, ADTree 1. The dataset used is not publicly available
2. An unknown number of records in the generated dataset
Ali Bin et al. (2022) [100] Kali Linux distribution having 3855 attacks evaluation of machine learning classifiers for the detection of man-in-the-middle attacks Decision Tree
Naive Bayes
K-Nearest Neighbor(KNN)
Artificial Neural Network
XGBoost
2. The result obtained is limited to only one dataset
Muhanna Saed and Ahamed Aljuhani (2022) [91] Wi-Fi network benchmark dataset Proposal machine learning techniques both to detect and identify Man-In-The-Middle attacks on a wireless communication network LSTM, Support Vector Machine, Random Forest 1. proposed model likely to be bias due to imbalance dataset
Kamal Omari (2023) [75] UCI phishing domains dataset an investigation into the performances and efficiency of Logistic Regression, KNN, SVM, Naive Bayes, Decision Tree, Random Forest, and Gradient Boosting for phishing detection task naive bayes
KNN
SVM
Random Forest
Gradient Boost
Decision Tree
Logistic Regression
Prince Roy et al. (2022) [100] Kaggle SQL Injection attack dataset evaluation of machine learning classifiers for the detection of SQL Injection attacks Logistic Regression
Naive Bayes
XGBoost
Random Forest
Adaboost
Ann Zeky et al (2023) [66] Internally generated dataset proposal of extraction based naive Bayes robust model for phishing detection with emphasis on the combination of webpage content and URL feature analysis naive Bayes
URL analyzation
webpage content extraction
Patricia Iglesias et al. [51] Coco and ILSVR dataset proposal of deep learning based CNN model for detecting drive-by road attack Convolutional Neural Network limited to drive-by road attack detection in stego images
mahdi bahaghighat (2023) [21] Grega Vrbančič phishing dataset performance comparison of phishing detection method based on several six different algorithm naive bayes
KNN
SVM
Random Forest
Gradient Boost
Logistic Regression
Nishitha U et al (2023) [73] Kaggle Phishing Dataset performance comparison of machine learning and deep learning based algorithm for phishing detection CNN
RNN
KNN
Random Forest
Decision Tree
Logistic Regression
2. imbalance in the dataset will lead to bias
sole reliance on URL feature
Prince Roy et al. (2022) [100] Kaggle SQL Injection attack dataset evaluation of machine learning classifiers for the detection of SQL Injection attacks Logistic Regression
Naive Bayes
XGBoost
Random Forest
Adaboost
Twana Mustafa and Murat Karabatak (2023) [71] UCL Phishing Dataset Performance Comparison of different Bayesian classifier based on different Feature Selection Algorithm naive bayes
individual FS
forward FS
backward FS
Plus-I takeaway-r FS
AR1 FS
no hint on why Plus-I takeaway-r FS works better best for naive bayes
Morufu Olalere dataset was internally generated from dmoz proposal of Naive Bayes model for detecting SQL injection attack Naive Bayes SQL syntax distribution in the dataset was not known, It is difficult to know how the proposed model will perform on Kaggle SQL injection dataset
Jaya T et al (2023) [57] UK-2011, SpamBase, and Spam Assassin datasets usage of frequency weightage of the words for unsupervised clustering of mail into spam and ham messages naive bayes
random forest
logistic regression
random tree
LTSM
Santhosh Raminedi et al (2023) [85] Internally collected phishing dataset evaluation of several machine learning and deep learning based algorithms for phishing detection using URL features ANN
SVM
KNN and Naive Bayes
Random Forest
Decision Tree
Logistic Regression
2. sole reliance on URL feature
@@ II Background Study
II-A Category of Current State of the art Phishing Detection model
II-A1 Bayesian-Based-Classifier
@@
II-B Non-Bayesian Based Classifier
II-B1 Decision Tree
@@ II-B5 K-Nearest Neighbor (KNN)
II-B6 Support Vector Machine (SVM)
@@
II-C Deep Learning Based Classifier
II-C1 Convolutional Neural Network (CNN)
@@
II-C2 Recurrent Neural Network (RNN)
@@
II-C3 Long Short-Term Memory (LSTM)
@@ III Literature Review
TABLE II: Comparison Analysis of Methodologies for Detection of SQL Injection Attack
Classifier Authors Mean Score
Naive Bayes [89], [112], [30], [62], [78], [98], [9] 90.4
SVM [112], [30], [42], [62], [98], [45],[9] 87.63
Random Forest [89], [112], [98], [45], [105], [10], [9] 93.71
Logistic Regression [89], [62], [45], [18], [10], [9], [22] 89.68
KNN [98], [10], [9], [2], [43], [22], [54] 87.2
Decision Tree [45], [30], [105], [9], [2], [43], [52] 90.04
TABLE III: Comparison Analysis of Methodologies for Detection of DDoS Attack
Classifier Authors Mean Score
SVM [102], [64], [44], [16], [79] 90.0
Naive Bayes [102], [93], [44], [5], [88], [116] 84.6
Random Forest [8], [93], [44], [36], [5], [88] 93.34
Decision Tress [8], [102], [44], [36], [16], [106] 96
XGBoost [8], [36], [70], [25], [90] 96.2
KNN [102], [36], [16], [32], [115] 96.5
TABLE IV: Comparison Analysis of Methodologies for Detection of Phishing Attack
Classifier Authors Mean Score
Naive Bayes [21], [85], [114], [58], [75], [6] 80.431
SVM [21], [85], [58], [75], [14], [6] 89.429
Random Forest [21], [85], [114], [58], [75], [13] 97.065
Decision Tree [85], [58], [75], [14], [6], [86] 95.248
Logistic Regression [21], [85], [75], [17], [76], [101] 92.589
KNN [21], [85], [58], [75], [6], [11] 90.479
TABLE V: Comparison Analysis of Methodologies for Malware Detection
Classifier Authors Mean Score
SVM [72], [107], [67], [35], [59], [97] 85.5
Naive Bayes [67], [69],[94], [65], [96], [35] 79.6
Random Forest [67], [80], [28], [95], [3], [69] 96.04
Decision Tree [67], [95], [3], [69], [65], [97] 91.6
KNN [72], [65], [67], [96], [35], [97] 86.6
Logistic Regression [65], [96], [20], [99], [69], [63] 90.06
TABLE VI: Comparison Analysis of Methodologies for Man-in-the-Middle Attack Detection
Classifier Authors Mean Score
SVM [29], [68], [15], [39], [53], [41] 92.3
Naive Bayes [68], [7], [110], [61], [108], [82] 82.8
Random Forest [29], [7], [110], [41], [31], [87] 96.6
Decision Tree [29], [68], [15], [41], [31], [61] 88.6
KNN [29], [68], [7], [15], [110], [61] 91.5
Logistic Regression [29], [41], [61], [82], [104], [24] 84.8
TABLE VII: Comparison Analysis of Methodologies for Drive-By Attack Detection
Classifier Authors Mean Score
SVM [1], [81], [26], [12], [83], [55] 88.69
Naive Bayes [56], [1], [47], [103], [77], [92] 87.37
Random Forest [1], [12], [83], [55], [77], [23] 91.83
Decision Tree [1], [40], [60], [55], [103], [34] 91.85
KNN [81], [55], [103], [77], [92], [74] 92.22
Logistic Regression [81], [55], [77], [34], [27], [109] 92.76
IV Analysis and Discussion
@@ 3. Limitation of Current Approach to SQL Injection Attack Detection
V Conclusion and Research Direction